# Closed-form finite-time blow-up and conditional perturbative stability framework for a (1+2)(1+2)D system (E1) derived from the 2D inviscid Boussinesq equations

### 31 Mar 2026 v2

@@ 3. Perturbation PDEs
We derive the remainder system around the explicit background. The full solutions are written as the sum of the background fields and the remainders:
@@
After substituting  (3.1) into  (2.6)(1,2,3,4), we separate the full system into the background equations for $(V,U,G,P)$ and the exact remainder equations for $(v,u,g,p)$. In the present revision all pure-background terms are kept in the background equations, so the remainder system contains only linear couplings to the background and genuinely nonlinear remainder–remainder interactions:
$\left\{\begin{aligned} \lambda u_{t}&=\tfrac{1}{2}(uV+Uv)-(g+v)\xi U_{\xi}-(G+V)\xi u_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xu_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xU_{x}+N_{1}\\[5.69054pt] \lambda v_{t}&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\qquad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{1}{\xi}p_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x}+N_{2}+\tfrac{1}{x}p_{x}\\[5.69054pt] \lambda g_{t}&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\quad\quad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{\mu}{\xi}p_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}+N_{3}+\tfrac{\mu}{x}p_{x}\\[5.69054pt] 0&=\xi xg_{x}+xv_{x}-({\xi^{2}+1})(g-v+\xi g_{\xi}+\xi v_{\xi})\end{aligned}\right$. (3.2)
@@ 3.1. Getting rid of $(p_{\xi},p_{x})$
Our next step is to get rid of $(p_{\xi},p_{x})$. Define $\omega$ as
@@
$\left\{\begin{aligned} M_{2}&=\tfrac{10}{3\xi}\left(\left(\xi^{2}+1\right)u_{\xi}+\xi xu_{x}\right)u\\ &+\tfrac{\xi^{2}-1}{\xi^{2}+1}\left(x\psi_{x}+2\psi\right)\omega\\ &-\xi\left(x\psi_{x}+2\psi\right)\omega_{\xi}\\ &+\big(\xi x\psi_{\xi}-\tfrac{\xi^{2}-1}{\xi^{2}+1}x\psi\bigr)\omega_{x}.\end{aligned}\right$. (3.12)
@@ 4.1. Background Solutions
The background solutions $V(t,x,\theta),U(t,x,\theta)$ are updated with the new seed profiles from (4.2):
@@
Initial conditions
$r:=x^{2}\bigl(1+\cos^{2}(2\theta)\bigr)\geq 0,\qquad V_{0}(x,\theta)=\frac{A}{1+r^{3}},\qquad U_{0}(x,\theta)=\frac{B\,r}{\bigl(1+r^{3}\bigr)^{2}}.$ (4.2)
@@ 4.2. Initial energy and finiteness
$E(0):=\int_{-\pi/4}^{\pi/4}\int_{-\infty}^{\infty}x^{2}\Big[(V_{0}(x,\theta))^{2}+U_{0}^{2}(x,\theta)\Big]\ |x|\,dx\,d\theta.$ (4.3)
Remark 4.1 (Behavior of the weighted energy for the new seed choice).
In the modified background model (with $U^{2}$ in the VV–equation) we take the energy density $V^{2}+U^{2}$. For the new seeds (4.2),
$r:=x^{2}\bigl(1+\cos^{2}(2\theta)\bigr),\qquad V_{0}(x,\theta)=\frac{A}{1+r^{3}},\qquad U_{0}(x,\theta)=\frac{B\,r}{(1+r^{3})^{2}},$
one still has $V_{0}(0,\theta)=A$ and $U_{0}(0,\theta)=0$ for every $\theta\in[-\pi/4,\pi/4]$, so the blow-up time remains $T=6/A$ and the singularity remains localized at the origin $x=0$. A Taylor expansion at $r=0$ gives
$V_{0}(x,\theta)=A\bigl(1-r^{3}+O(r^{6})\bigr),\qquad U_{0}(x,\theta)=B\,r+O(r^{4}),$
while at spatial infinity one has $r\sim x^{2}$, hence
$V_{0}(x,\theta)=O(x^{-6}),\qquad U_{0}(x,\theta)=O(x^{-10}).$
Therefore the initial weighted energy $E(0)$ is finite. The endpoint behavior at $t=T$ is different, however: the new seeds keep the background explicit and smooth for every fixed $t<T$, but they no longer produce a uniformly bounded endpoint energy at $t=T$.
@@ 5. Linear estimates and conditional nonlinear control up to blow-up time
5.1. Bootstrap framework and updated background coefficient bounds
@@
Background coefficient bounds actually needed in the energy method. The explicit background (4.1) provides $(U,V)$ in closed form. What the stability estimates require is not a uniform bound on the raw derivatives $\partial_{x}^{m}\partial_{\theta}^{\ell}V$ (which can grow faster than $(T-t)^{-1}$ near the intermediate scale $r^{2}\sim T-t$), but rather uniform control of the degenerate combinations that appear in (3.14),(3.11) and in the weighted Sobolev norms.
@@
The estimate is a direct consequence of the closed-form background formula together with the seed expansions near the ridge apex and the fact that away from the apex the denominator in (4.1) is bounded below uniformly. Near the apex one uses the denominator scale $D(t,x,\theta)\gtrsim(T-t)^{2}+r^{2}$, together with the observation that $Z_{x}=x\partial_{x}$ acts like $r\partial_{r}$ and $D_{\xi}$ neutralizes the harmless trigonometric singular factors in angular derivatives. ∎
@@ 5.3. Energy inequality for $(u,\omega)$
Assume that the explicit background solution exists on $[0,T)$, blows up at $(x,\xi)=(0,\pm 1)$ at time $T=6/A$, and satisfies the adapted coefficient bounds of Lemma 5.1. Assume also that the weighted elliptic estimate (5.3) holds. Let $k\geq 6$, and let $(u,\omega,\psi)$ solve the exact remainder system on $0,t_{*}]\subset[0,T)$.
@@
Accordingly, the logical bottleneck of the manuscript is no longer a forcing obstruction in the remainder equations. The main unresolved issue is instead the rigorous construction of the background solution up to its blow-up time with the required ridge flatness and coefficient bounds, together with whatever refined estimate is needed to produce a genuine gap exponent $\sigma<1$ in the remainder norm. Once those two inputs are available, the present stability mechanism upgrades directly to a nonlinear remainder theorem in the spirit of Elgindi.
@@ 6. Conclusion
We derived a closed (1+2)(1+2)D subsystem (E1) from the inviscid 2D Boussinesq equations under a parity ansatz and organized its blow-up analysis around two components: a singular background core and an exact remainder system. In the updated formulation, the ridge rays carry a convection-free (1+1)(1+1)D CLM-type reaction dynamics with explicit finite-time blow-up, while the full wedge problem is rewritten in the $(x,\xi)$ variables so that the remainder equations are exact and all pure-background terms remain in the background system.
The main gain of this revision is conceptual and technical at the same time. On the remainder side, the earlier large forcing obstruction has disappeared. What remains is a singular linear problem with quadratic lower-order interactions, and this is precisely the setting in which an Elgindi-type transfer mechanism becomes available. The weighted energy method developed in Section 5 now shows that, if the background exists on $[0,T)$ with the coefficient bounds required there, and if the remainder stays subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
Accordingly, the main unresolved step is now sharply identified. It is no longer the derivation of the remainder system, nor the linear weighted estimate, but the full-wedge background construction up to the blow-up time together with the rigidity properties needed to propagate the ridge structure and close the nonlinear bootstrap without loss. Revision (c) makes this explicit: the ridge blow-up mechanism itself is available in closed form, but extending it to a full background with the necessary compatibility bounds remains the decisive open problem.
This sharpened formulation is useful even before the final nonlinear theorem is complete. It separates what is already rigorous from what is still conditional, and it places the manuscript much closer in spirit to the singular-core-plus-remainder logic used by Elgindi and collaborators. At the present stage, the paper provides an exact derivation from Boussinesq, an explicit ridge blow-up mechanism, an exact remainder system, singular weighted linear control, and a precise conditional blow-up transfer statement.
Natural next steps are therefore clear. The first is to prove the full background existence and blow-up theorem for the anisotropic seed family used here. The second is to sharpen the detecting norm so that the remainder remains strictly below the background blow-up rate, yielding a closed nonlinear bootstrap. After that, one can revisit modulation of geometric parameters and lower-regularity weighted theories.
@@ 7. Acknowledgements
ChatGPT is credited as a substantive contributor to drafting and technical editing; responsibility for correctness remains with the author. The author thanks Prof. Zixiang Zhou of the Department of Mathematics at Fudan University and Prof. Jie Qin of the Department of Mathematics at the University of California, Santa Cruz for their continuous support and encouragement over the years. The author also thanks colleagues and the broader PDE and fluid-dynamics community for stimulating discussions on inviscid Boussinesq and CLM-type models. (Computational assistance: ChatGPT, GPT–5.2 Thinking, OpenAI; sessions in March 2026.)

### 03 Apr 2026 v3

@@ 3. Perturbation PDEs
We derive the remainder system around a prescribed background whose ridge/apex behavior is the one identified above. The full solutions are written as the sum of the background fields and the remainders:
@@
After substituting  (3.1) into  (2.8)(1,2,3,4), we separate the full system into the background equations for $(V,U,G,P)$ and the exact remainder equations for $(v,u,g,p)$. All pure-background terms are kept in the background equations, so the remainder system contains only linear couplings to the background and genuinely nonlinear remainder–remainder interactions:
$\left\{\begin{aligned} \lambda u_{t}&=\tfrac{1}{2}(uV+Uv)-(g+v)\xi U_{\xi}-(G+V)\xi u_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xu_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xU_{x}+N_{1}\\[5.69054pt] \lambda v_{t}&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\qquad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\xi p_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x}+N_{2}+\tfrac{1}{x}p_{x}\\[5.69054pt] \lambda g_{t}&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\quad\quad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{\mu}{\xi}p_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}+N_{3}+\tfrac{\mu}{x}p_{x}\\[5.69054pt] 0&=\xi xg_{x}+xv_{x}-({\xi^{2}+1})(g-v+\xi g_{\xi}+\xi v_{\xi})\end{aligned}\right$. (3.2)
@@ 3.1. Getting rid of $(p_{\xi},p_{x})$
Our next step is to get rid of $(p_{\xi},p_{x})$. Define $(\omega,\Omega)$ as
@@
$\left\{\begin{aligned} M_{2}&=\tfrac{10}{3\xi}\left(\left(\xi^{2}+1\right)u_{\xi}+\xi xu_{x}\right)u\\ &+\tfrac{\xi^{2}-1}{\xi^{2}+1}\left(x\psi_{x}+2\psi\right)\omega\\ &-\left(x\psi_{x}+2\psi\right)\xi\omega_{\xi}\\ &+\big(\xi\psi_{\xi}+\tfrac{1-\xi^{2}}{\xi^{2}+1}\psi\bigr)x\omega_{x}.\end{aligned}\right$. (3.12)
@@ 4.1. Background Solutions
The formulas below define a candidate background pair $V(t,x,\xi),U(t,x,\xi)$ associated with the end-vanishing seed profiles from (2.23). They agree with the explicit ridge reduction at the apex $(x,\xi)=(0,\pm 1)$ and are used throughout as the background ansatz for the conditional perturbative analysis:
@@
The initial seed profiles are
$\left\{\begin{aligned} &r=x^{2}+A_{1}\phi(\xi),\\ &a(x,\xi)=\frac{A}{(1+r)^{3}},\\ &b(x,\xi)=\frac{B\,r}{(1+r)^{6}},\end{aligned}\right$. (4.2)
with $\phi$ given by (2.21). Along the ridge endpoints $\xi=\pm 1$ one has $\phi(\pm 1)=0$, hence
$r\big|_{\xi=\pm 1}=x^{2},\qquad a(x,\pm 1)=\frac{A}{(1+x^{2})^{3}},\qquad b(x,\pm 1)=\frac{B\,x^{2}}{(1+x^{2})^{6}},$
which recovers the explicit apex/ridge reduction used earlier.
Remark 4.1 (Series expansion of the new seed near the apex points).
Fix an endpoint $\xi_{0}\in\{\pm 1\}$, and write $\delta:=\xi-\xi_{0}$. Since
$\phi(\xi)=\exp\bigl(-(\xi^{2}-1)^{-2}\bigr)$
is flat at $\xi=\xi_{0}$, one has
$\phi(\xi)=O(|\delta|^{N})\qquad\text{for every }N\in\mathbb{N}$
as $\xi\to\xi_{0}$. Hence near $(x,\xi)=(0,\xi_{0})$,
$r=x^{2}+A_{1}\phi(\xi)=x^{2}+O(|\delta|^{N})\qquad\forall N\in\mathbb{N}.$
Therefore the seed profiles admit the local expansions
$a(x,\xi)=\frac{A}{(1+r)^{3}}=A-3Ar+6Ar^{2}+O(r^{3}),$
$b(x,\xi)=\frac{Br}{(1+r)^{6}}=Br-6Br^{2}+21Br^{3}+O(r^{4}),$
that is,
$a(x,\xi)=A-3Ax^{2}+O\bigl(x^{4}+\phi(\xi)\bigr),$
$b(x,\xi)=Bx^{2}+A_{1}B\phi(\xi)+O\bigl(x^{4}+x^{2}\phi(\xi)+\phi(\xi)^{2}\bigr)$
near $(0,\xi_{0})$. In particular,
$a(0,\xi_{0})=A,\qquad b(0,\xi_{0})=0,\qquad\partial_{\xi}^{k}a(0,\xi_{0})=\partial_{\xi}^{k}b(0,\xi_{0})=0\quad(k\geq 1).$
Thus the new seed is smooth and non-singular at the apex points, and its angular dependence is super-flat there. The only non-analytic feature is the standard flatness of $\phi$ at the endpoints, which is compatible with the apex reduction used in this manuscript. By contrast, the alternative seed
$r=x^{2}\bigl(1+A_{1}\phi(\xi)\bigr)$
would force $r=0$ at the origin $x=0$, so the present choice retains more transverse structure away from the apex while still giving the same leading jet at $(0,\pm 1)$.
Remark 4.2 (Seed 7 versus Seed 8 near the apex).
For the current seed choice
$\text{Seed 7:}\qquad r=x^{2}+A_{1}\phi(\xi),$
the top view near the apex displays a clear and pronounced “++” pattern centered at the apex point. For the alternative choice
$\text{Seed 8:}\qquad r=x^{2}\bigl(1+A_{1}\phi(\xi)\bigr),$
the same “++” pattern is smeared near the apex point because the angular modulation is multiplied by the vanishing factor $x^{2}$. In particular, Seed 7 preserves a sharper apex profile, while Seed 8 gives a more blended local geometry.
For the purposes of the present paper, however, both Seed 7 and Seed 8 are admissible. They produce the same apex values and the same leading-order ridge/apex jet needed for the pointwise reduction at $(x,\xi)=(0,\pm 1)$. Thus either seed can be used within the current framework, although Seed 7 gives a visually clearer apex structure.
@@ 4.2. Initial energy and finiteness
$E(0):=\int_{-1}^{1}\int_{-\infty}^{\infty}x^{2}\Big[a(x,\xi)^{2}+b(x,\xi)^{2}\Big]\ \frac{|x|}{1+\xi^{2}}\,dx\,d\xi.$ (4.3)
Remark 4.3 (Behavior of the weighted energy for the end-vanishing seed choice).
For the background ansatz (4.1) built from (2.23), one has
$r=x^{2}+A_{1}\phi(\xi),\qquad a(x,\xi)=\frac{A}{(1+r)^{3}},\qquad b(x,\xi)=\frac{B\,r}{(1+r)^{6}}.$
Since $0\leq\phi(\xi)\leq 1$ on [1,1][-1,1], one has $r\sim 1+x^{2}$ uniformly in $\xi$, and therefore
$a(x,\xi)=O\bigl((1+x^{2})^{-3}\bigr),\qquad b(x,\xi)=O\bigl((1+x^{2})^{-5}\bigr)\quad\text{as }|x|\to\infty,$
uniformly for $\xi\in[-1,1]$. Hence the weighted initial energy (4.3) is finite.
At the ridge endpoints one has $\phi(\pm 1)=0$, so
$a(0,\pm 1)=A,\qquad b(0,\pm 1)=0,$
and the blow-up time induced by the ridge ODE remains
$T=\frac{6}{A}.$
More generally, every $\xi$-derivative of $\phi$ vanishes at $\xi=\pm 1$, so all endpoint angular jets of $a$ and $b$ vanish there; this is the main reason the new seed family is compatible with the apex-only ridge analysis developed in the present paper.
4.3. Initial pressure determined by the PPE
At $t=0$, the explicit background satisfies
$U(0,x,\xi)=b(x,\xi)=\frac{B\,r}{(1+r)^{6}},\qquad r=x^{2}+A_{1}\phi(\xi).$
Therefore
$U(0,x,\xi)^{2}=\frac{B^{2}r^{2}}{(1+r)^{12}}=:F(r),\qquad F^{\prime}(r)=\frac{2B^{2}r(1-5r)}{(1+r)^{13}}.$
Since
$r_{x}=2x,\qquad r_{\xi}=A_{1}\phi^{\prime}(\xi),$
one obtains
$(U(0)^{2})_{x}=2x\,F^{\prime}(r),\qquad(U(0)^{2})_{\xi}=A_{1}\phi^{\prime}(\xi)F^{\prime}(r).$
Substituting these formulas into the Poisson pressure equation (2.12) yields the initial elliptic problem for
$P_{0}(x,\xi):=P(0,x,\xi):$
$P_{0,xx}+\frac{1}{x}P_{0,x}+\frac{2(1+\xi^{2})\xi}{x^{2}}P_{0,\xi}+\frac{(1+\xi^{2})^{2}}{x^{2}}P_{0,\xi\xi}$ (4.4)
$\qquad=\frac{B^{2}r^{2}}{(1+r)^{12}}-\frac{4A_{1}B^{2}}{x}\,(1+\xi^{2})^{1/2}\,\xi\,\phi^{\prime}(\xi)\,\frac{r(1-5r)}{(1+r)^{13}}+4B^{2}x(1+\xi^{2})^{-1/2}\,\frac{r(1-5r)}{(1+r)^{13}},$
where $r=x^{2}+A_{1}\phi(\xi)$. Thus the initial pressure is not prescribed independently: it is determined by the elliptic solve (4.4), together with the symmetry/decay conditions imposed on $P$ in the wedge formulation.
@@ 5. Linear estimates and conditional nonlinear control up to blow-up time
5.1. Bootstrap framework and adapted background coefficient bounds
@@
Background coefficient bounds actually needed in the energy method. The candidate background (4.1) provides a closed-form model for $(U,V)$ and, in particular, reproduces the explicit apex dynamics. What the stability estimates require is not a uniform bound on the raw derivatives $\partial_{x}^{m}\partial_{\theta}^{\ell}V$ (which can grow faster than $(T-t)^{-1}$ near the intermediate scale $r^{2}\sim T-t$), but rather uniform control of the degenerate combinations that appear in (3.14),(3.11),(3.12) and in the weighted Sobolev norms.
@@
The estimate is a direct consequence of the closed-form background formula together with the seed expansions near the ridge apex and the fact that away from the apex the denominator in (4.1) is bounded below uniformly. Near the apex one uses the denominator scale $D(t,x,\xi)\gtrsim(T-t)^{2}+r^{2}$, together with the observation that $Z_{x}=x\partial_{x}$ acts like $r\partial_{r}$ and $D_{\xi}$ neutralizes the harmless trigonometric singular factors in angular derivatives. ∎
@@ 5.3. Energy inequality for $(u,\omega)$
Assume that a compatible background solution exists on $[0,T)$, has the same apex blow-up rate as the explicit ridge dynamics at $(x,\xi)=(0,\pm 1)$ with time $T=6/A$, and satisfies the adapted coefficient bounds of Lemma 5.1. Assume also that the weighted elliptic estimate (5.3) holds. Let $k\geq 6$, and let $(u,\omega,\psi)$ solve the exact remainder system on $0,t_{*}]\subset[0,T)$.
@@
Accordingly, the logical bottleneck of the manuscript is no longer a forcing obstruction in the remainder equations. The main unresolved issue is instead the rigorous construction/control of a background away from the apex, with the coefficient bounds needed by the weighted energy method and with enough rigidity near the apex to match the explicit ridge dynamics, together with whatever refined estimate is needed to produce a genuine gap exponent $\sigma<1$ in the remainder norm. Once those two inputs are available, the present stability mechanism upgrades directly to a nonlinear remainder theorem in the spirit of Elgindi.
@@ 6. Conclusion
We derived a closed (1+2)(1+2)D subsystem (E1) from the inviscid 2D Boussinesq equations under a parity ansatz and organized its blow-up analysis around two components: an explicit ridge/apex core and an exact remainder system. In this formulation, the ridge rays carry a convection-free (1+1)(1+1)D CLM-type reaction dynamics with explicit finite-time blow-up at the apex $(x,\xi)=(0,\pm 1)$, while the full wedge problem is rewritten in the $(x,\xi)$ variables so that the remainder equations are exact and all pure-background terms remain in the background system.
The weighted energy method developed in Section 5 shows that, if a compatible background exists on $[0,T)$ with the coefficient bounds required there and with apex trace governed by the explicit ridge dynamics, and if the remainder stays subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
The main unresolved step is therefore the construction and control of a full background away from the apex, together with the rigidity properties needed to match the apex dynamics and close the nonlinear bootstrap without loss. The blow-up mechanism itself is available in closed form at the apex, but extending that information to a full background with the necessary compatibility bounds remains the decisive open problem.
Even before the final nonlinear theorem is completed, the present formulation already isolates the core components of the analysis. It provides an exact derivation from Boussinesq, an explicit apex blow-up mechanism, a strong linearized stability framework, conditional nonlinear control, and a precise conditional blow-up transfer statement.
Natural next steps are therefore clear. The first is to prove the full background existence/control theorem compatible with the apex dynamics identified here. The second is to sharpen the detecting norm so that the remainder remains strictly below the background blow-up rate, yielding a closed nonlinear bootstrap. After that, one can revisit modulation of geometric parameters and lower-regularity weighted theories.
@@ 7. Acknowledgements
ChatGPT is credited as a substantive contributor to drafting and technical editing; responsibility for correctness remains with the author. The author thanks Prof. Zixiang Zhou of the Department of Mathematics at Fudan University and Prof. Jie Qin of the Department of Mathematics at the University of California, Santa Cruz for their continuous support and encouragement over the years. The author also thanks colleagues and the broader PDE and fluid-dynamics community for stimulating discussions on inviscid Boussinesq and CLM-type models. (Computational assistance: ChatGPT, GPT–5.2 Thinking,GPT–5.4 Thinking, OpenAI; sessions in March 2026.)